OpenStudy (anonymous):

if a pizza with a diameter of 12 inches costs 10 99$ based on area about how much should a 15-inch pizza costs?

OpenStudy (mertsj):

\[\frac{12}{10.99}=\frac{15}{x}\]

OpenStudy (anonymous):

that's not right. I got 13.7 out of that..

OpenStudy (mertsj):

Or based on area: \[\frac{36\pi}{10.99}=\frac{7.5^2\pi}{x}\]

OpenStudy (anonymous):

it's based on area.

OpenStudy (mertsj):

17.17

OpenStudy (anonymous):

thanks!

OpenStudy (ganpat):

yup 17.17 $, sorry i missed that..